OpenStudy (anonymous):

A cable that weighs 4 lb/ft is used to lift 850 lb of coal up a mine shaft 700 ft deep. Find the work done. (looking just for the Riemann sum and Integral setup, I'll find the answer no problem from there, and will give a medal for doing so)

OpenStudy (anonymous):

\[\int\limits_{0}^{700}(4x) dx \] is incorrect

OpenStudy (anonymous):

Clearly missing something, and X\(_i\) isn't taking as... (writes)

OpenStudy (anonymous):

cable weight is 4*700

OpenStudy (anonymous):

so total weight is 4*700 +850

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \sum_{i=1}^{n} (4)(x_i*) \triangle x\]

OpenStudy (anonymous):

Is there something wrong with my sum there?

OpenStudy (mathsolver):

can i offer an alternative solution? if you take the average weight of the cable. i.e 1400. and then add that to the 850, you will end up with the average weight you have to lift. I would then work out the work done from there

OpenStudy (anonymous):

should be probably 4*(700-x)

OpenStudy (anonymous):

\[Work = \int\limits_{a}^{b} force\cdot distance\] Correct?

OpenStudy (anonymous):

Well and cos \(\theta\), but we don't have that for this problem

OpenStudy (anonymous):

\[\int\limits_{0}^{700}4(700-x) + 850 dx\]

OpenStudy (mathsolver):

\[\int\limits_{0}^{700}(850-4x)dx \] I think this is your final answer

OpenStudy (anonymous):

850 is the weight that is there till the end. The weight of cable is changing.

OpenStudy (anonymous):

you are resting the cable weight from the coal weight...?

OpenStudy (anonymous):

i think my setup is correct, just express lb in newtons

OpenStudy (anonymous):

1 lb-f ≈ 4.45 N

OpenStudy (anonymous):

so should be: \[4,45\int\limits_{0}^{700}[4*(700-x)+850]dx\]

OpenStudy (mathsolver):

I agree @myko, however there is a much easier way to solve this

OpenStudy (anonymous):

Got it! You were dead on mathsolver and much to my disbelief it was because I had this: \[\lim_{n \rightarrow \infty} \sum_{i=1}^{n} (4)(x_i*) \triangle x - 850*700\]

OpenStudy (anonymous):

It's supposed to be added >_<

OpenStudy (anonymous):

And now it's correct, what do you know lol

OpenStudy (anonymous):

Thank you all for helping!